You plan to invest ​$115,767.00, borrowing 68% of the money ($78,721.56) from a​ friend, so that you will put up ​$37045.44 of your own money. The cost of debt​ (interest rate on the​ loan) is 11.16% and there are no taxes or transaction costs. With the​ arrangement, you expect a return of 18.6% on your equity investment. What would be your expected rate of return without the​ debt? That​ is, what rate of return would you expect to earn on your money if the entire ​$115,767.00 was your own​ money? ​(Percent with 2​ decimals.)
Answer #1

Return on Investment = 0.186%

Return = 0.186(37,045.44) = $6,890.45

Cost of Debt = 0.1116(78,721.56) = $8,785.33

Expected Rate without Debt = (6,890.45 + 8,785.33)/115,767

Expected Rate without Debt = 13.54%